OpenStudy (anonymous):

Three Balls (A,B,C) all have the same mass and are all dropped from the same height. Ball A rebounds odd the floor the fastest, Ball C comes to a stop on the floor and doesn't rebound at all. They all have the same initial speed right before contact with the floor. Ball B rebounds a little, but not much. Which ball has the largest change in momentum? Which ball has the smallest change in momentum? Which ball experiences the largest force from the ground? Which ball experiences the largest acceleration? The answers can be A,B,C, or all same

OpenStudy (anonymous):

Well according to Galileo they all have the same acceleration until impact g=9,91m/s^2. So What does Isaac Newton tell us? He says F=ma and.... so on. So the ball who rebounce the most must have retained most force am I right?

OpenStudy (anonymous):

You question has no meaning in context lol.

OpenStudy (anonymous):

\[\Delta P= m \Delta v\] For the three balls: \(\Delta_A=u-(-u)=2u \) \(\Delta_B=v-(-u)=u+v, where, v<u \) \(\Delta_C=0-(-u)=u \) thus A has largest change in momentum, C smallest. \(F=\frac{dP}{dt}\) The larger the change in momentum wrt time, the larger the force applied. Thus A has the largest force applied, and \(F=ma \) so larger force produce larger acceleration. Thus A has the largest acceleration.